What is the solution of Y and how did you get it?
6 = -y + 42 - 2y
THANKS!!!!!!!!!

Answer by Alan3354(69443)

6 = -y + 42 - 2y
Combine -y & -2y
6 = 42 - 3y
Add 3y
3y + 6 = 42
Subtract 6
3y = 36
Divide by 3
y = 12